OpenStudy (anonymous):

Find all real values of x, such that f(x)=0. f(x)=15-3x

OpenStudy (anonymous):

That only means that you have to solve f(x)=0 15-3x=0 15=3x 5=x

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

what about find the value(s) of x for which f(x)=g(x), given f(x)=x^2 and g(x)=x+2

OpenStudy (anonymous):

you can set them equal to each other and then solve for x

OpenStudy (anonymous):

Than you have to solve: x^2=x+2 x^2-x-2=0 (x-2)(x+1)=0 x=2 or x=-1

OpenStudy (anonymous):

thank you so much!